Read the question carefully. Sample question: “80% of people who purchase pet insurance are women. If 9 pet insurance owners are randomly selected, find the probability that exactly 6 are women.”
I am trying to have this problem broken down. Can you please explain how 84 came from 9! / (9-6)!6!

Found 2 solutions by Theo, addingup:
Answer by Theo(13342)

p(x) = p^x * q^(n-x) * c(n,x)

p(x) is the probability of exactly occurrences.
n = number of total occurrences.
p = probability of success on one trial.
q = probability of failure on one trial.
c(n,x) = number of ways you can get a set of 8 out of a set of n.

n = 9
x = 6
p = .8
q = 1 - .8 = .2
c(9,6) = 9! / (6! * 3!) = 84

p(6) = .8^6 * .2^3 * 84 = .176160768

the sum of all probabilities must be equal to 1.

the following excel spreadsheet shows this to be true.

look below the spreadsheet for the answer to your question.

$$$

Can you please explain how 84 came from 9! / (9-6)!6!

c(n,x) = n! / (x! * (n-x)!)

in your problem, this becomes c(9,6) = 9! / (6! * (9-6)!)

simplify this to get c(9,6) = 9! / (6! * 3!)

simplify further to get c(9,6) = (9*8*7*6*5*4*3*2*1) / (6*5*4*3*2*1 * 3*2*1)

you have a block of (6*5*4*3*2*1) in the numerator and the denominator that cancels out and you are left with c(9,6) = (9*8*7) / (3*2*1)

the 3 in the denominator goes into the 9 in the numerator 3 times.
the 2 in the denominator goes into the 8 in the numerator 4 times.

you are left with c(9,6) = 3*4*7 which is equal to 12 * 7 which is equal to 84.











Answer by addingup(3677)

9! = 1 x 2 x 3 x 4 x 5 x 6 x 7 x 8 x 9
(9-6)! = 3! = 1 x 2 x 3
6! = 1 x 2 x 3 x 4 x 5 x 6
----------------------------
A factorial (indicated by the sign !) means the product of all the whole numbers from 1 to n, so it's 1 x 2 x 3 x .... n
Look at your calculator and see if it has a factorial (!) key (mine does). If not, find the factorials then do the rest of the operations and if you do them right you will get 84.
NOTE: You don't have to do 1 times 2, skip the 1. Any number multiplied by 1 = the number, so there's no point doing 1 x 2 = 2. Just begin with 2 x 3.